# Beyond MP2 initialization for unitary coupled cluster quantum circuits

### v3

@@ Abstract
Despite its current popularity, VQE possesses a number of drawbacks. In particular, the classical optimization of circuit parameters presents many challenges, including barren plateaus (i.e., exponentially vanishing gradients in high dimensions), local minima, and saddle points [51, 52, 53, 54, 55]. Many approaches exist for minimizing the computational burden of classical optimization for VQE, with some proposals eschewing optimization entirely [56, 57, 50, 58, 59]. The crux of several of these strategies is a focus on choosing high-quality initial parameters, shifting some of the computational burden from optimization to initialization. While taking this idea to the limit, in which classical computation can produce a circuit that generates a wave function beyond what we can simulate classically seems unlikely with current technologies, pushing this limit as far as can be done with high performance computing resources is likely to be a fruitful strategy for near term and fault tolerant eras. In this work we compare the utility of different initialization strategies for a particular VQE ansatz that is often employed in quantum chemistry problems, the unitary coupled cluster (UCC) ansatz [60, 61, 62, 63, 64, 65].
There are several proposed strategies for generating the initial parameters for the UCC ansatz [66, 67, 64, 68, 69], including applications in which no optimization is performed on quantum hardware [58]. The most widely employed strategy generates parameters using classical Møller–Plesset perturbation theory of the second order (MP2) [66, 67, 64, 68]. Another less thoroughly studied approach is the use of the coupled cluster singles and doubles (CCSD) classical simulation method to generate initial parameters [70, 71]. The CCSD technique generally produces more accurate ground state energies than MP2 calculations and is not prohibitively more computationally burdensome for all but the largest of problems. However, CCSD has not been previously explored as a method to initialize VQE circuits, partly because the sparse wave function simulators in which it can be tested were only recently developed, and partly because it often produces non-variational energies for strongly correlated molecules. This raises the question: Which technique produces superior initial parameters for UCC ansatzes, MP2 or CCSD?
@@
Mol ASCI CCSD(T) CCSD
UCC
MP2
UCC
@@
It is well-known that traditional (projected) classical coupled cluster techniques can, in some situations, obtain energies that are not variational (dropping below the FCI results) or, even worse, energies that diverge from the physical ground state result. One such failure scenario is the chemistry of bond breaking, which we investigate here using the $\mathrm{H}_{10}$ molecule with a stretched interatomic distance of 1.5 Å. The CCSD and CCSD(T) energies of this molecule are lower than the FCI energy, an example of coupled cluster techniques producing non-variational results. Despite the failure of CCSD to produce an accurate energy for this molecule, the UCC circuit parameterized with CCSD produces a variational energy, as it must because the VQE approach is variational. The UCC (CCSD) energy for stretched $\mathrm{H}_{10}$ is 12.2% higher than the FCI energy, compared to 1.7% higher for the equilibrium geometry. These results show that the UCC ansatz parameterized with CCSD is robust to failures of the classical theory, but with some loss of accuracy. Regardless, our results show a close correspondence between UCC (CCSD) and CCSD theories and further study of this will provide insight into the power of coupled cluster approaches on quantum hardware.
@@
Finally, we briefly compare classical optimization of the UCC ansatz initialized with MP2 and CCSD parameters. As test cases we study LiH, $\mathrm{BeH}_{2}$, HF, and stretched $\mathrm{H}_{10}$, again using experimental geometries [78]. We consider both equilibrium and stretched geometries for LiH and $\mathrm{BeH}_{2}$, stretched by factors of 1.5 and 2.0 respectively, and the same stretched geometry of $\mathrm{H}_{10}$ as above. We use the cc-pCVDZ basis for LiH, the cc-pVDZ basis and the frozen core approximation for $\mathrm{BeH}_{2}$ and HF, and the STO-6G basis for stretched $\mathrm{H}_{10}$. We note that although the MP2 initialization sets $\theta=0$ for all singles operators, the circuits still contain these operators and their angles are included in the optimization. For each calculation we use $N_{\mathrm{WF}}=100,000$ and do not attempt to extrapolate energies to the infinite-$N_{\text{WF}}$ limit. Since the molecules in this study all have singlet spin states and geometries with symmetry, we can reduce the number of variational parameters that need to be optimized in the UCCSD ansatz in two ways. First, we use spin-complemented single- and double-excitation operators which combines parameters that are necessarily the same for these electronic states [75]. Second, we use point group symmetry arguments to eliminate parameters associated with operators that do not retain the symmetry of the initial Hartree-Fock SCF solution [82]. This reduces the number of parameters by a factor of approximately the order of the point group. For example, these symmetry considerations reduce the number of parameters in the LiH circuit from 2268 to 408. We use the BFGS method as implemented in SciPy to optimize the circuit parameters, setting the tolerance to $10^{-3}$, except for HF for which we set it to $5\times 10^{-3}$ due to limited computational resources [83]. We plot the errors as a function of optimization step in Fig. 5.

### v4

@@ Abstract
Despite its current popularity, VQE possesses a number of drawbacks. In particular, the classical optimization of circuit parameters presents many challenges, including barren plateaus (i.e., exponentially vanishing gradients in high dimensions), local minima, and saddle points [51, 52, 53, 54, 55]. Many approaches exist for minimizing the computational burden of classical optimization for VQE, with some proposals eschewing optimization entirely [56, 57, 50, 58, 59]. The crux of several of these strategies is a focus on choosing high-quality initial parameters, shifting some of the computational burden from optimization to initialization. While it seems unlikely with current technology for classical computation to produce a circuit capable of generating wave functions beyond what we can simulate classically, approaching this limit with high performance computing resources is likely to be a fruitful strategy for the near term and for the fault tolerant era. In this work we compare the utility of different initialization strategies for a particular VQE ansatz that is often employed in quantum chemistry problems, the unitary coupled cluster (UCC) ansatz [60, 61, 62, 63, 64, 65].
There are several proposed strategies for generating the initial parameters for the UCC ansatz [66, 67, 64, 68, 69], including applications in which no optimization is performed on quantum hardware [58]. The most widely employed strategy generates parameters using classical Møller–Plesset perturbation theory of the second order (MP2) [66, 67, 64, 68]. Another less thoroughly studied approach is the use of the coupled cluster singles and doubles (CCSD) classical simulation method to generate initial parameters [70, 71]. The CCSD technique generally produces more accurate ground state energies than MP2 calculations and is not prohibitively more computationally burdensome for all but the largest of problems. However, CCSD has not been previously explored as a method to initialize VQE circuits, partly because the sparse wave function simulators in which it can be tested were only recently developed, and partly because it often produces non-physical energies for strongly correlated molecules. This raises the question: Which technique produces superior initial parameters for UCC ansatzes, MP2 or CCSD?
@@
Mol ASCI CCSD(T) CCSD UCC (CCSD) MP2 UCC (MP2)
@@
It is well-known that traditional (projected) classical coupled cluster techniques can, in some situations, obtain energies that are not variational (dropping below the FCI results) or, even worse, energies that diverge from the physical ground state result. One such failure scenario is the chemistry of bond breaking, which we investigate here using the $\mathrm{H}_{10}$ molecule with a stretched interatomic distance of 1.5 Å. The CCSD and CCSD(T) energies of this molecule are lower than the FCI energy, an example of coupled cluster techniques producing non-variational results. Despite the failure of CCSD to produce an accurate energy for this molecule, the UCC circuit parameterized with CCSD produces a valid energy, as it must because the VQE approach is variational. The UCC (CCSD) energy for stretched $\mathrm{H}_{10}$ is 12.2% higher than the FCI energy, compared to 1.7% higher for the equilibrium geometry. These results show that the UCC ansatz parameterized with CCSD is robust to failures of the classical theory, but with some loss of accuracy. Regardless, our results show a close correspondence between UCC (CCSD) and CCSD theories and further study of this will provide insight into the power of coupled cluster approaches on quantum hardware.
Finally, we briefly compare classical optimization of the UCC ansatz initialized with MP2 and CCSD parameters. As test cases we study LiH, $\mathrm{BeH}_{2}$, HF, and stretched $\mathrm{H}_{10}$, again using experimental geometries [77]. We consider both equilibrium and stretched geometries for LiH and $\mathrm{BeH}_{2}$, stretched by factors of 1.5 and 2.0 respectively, and the same stretched geometry of $\mathrm{H}_{10}$ as above. We use the cc-pCVDZ basis for LiH, the cc-pVDZ basis and the frozen core approximation for $\mathrm{BeH}_{2}$ and HF, and the STO-6G basis for stretched $\mathrm{H}_{10}$. We note that although the MP2 initialization sets $\theta=0$ for all singles operators, the circuits still contain these operators and their angles are included in the optimization. For each calculation we use $N_{\mathrm{WF}}=100,000$ and do not attempt to extrapolate energies to the infinite-$N_{\text{WF}}$ limit. Since the molecules in this study all have singlet spin states and geometries with symmetry, we reduce the number of variational parameters that need to be optimized in the UCCSD ansatz in two ways. First, we use spin-complemented single- and double-excitation operators that combine parameters that are necessarily the same for these electronic states [75]. Second, we use point group symmetry arguments to eliminate parameters associated with operators that do not retain the symmetry of the initial Hartree-Fock SCF solution [79]. This reduces the number of parameters by a factor of approximately the order of the point group. For example, these symmetry considerations reduce the number of parameters in the LiH circuit from 2268 to 408. We use the BFGS method as implemented in SciPy to optimize the circuit parameters, setting the tolerance to $10^{-3}$, except for HF for which we set it to $5\times 10^{-3}$ due to limited computational resources [80]. We plot the errors as a function of optimization step in Fig. 5.